OpenStudy (anonymous):

Using substitution to solve the system x = 2y + 3 and 4x - 5y = 9, which of the following could be the first step?

OpenStudy (anonymous):

A. x - 2y = 3 4x - 5y = 9 B. 4(2y + 3) - 5y = 9 C. x + 2y = 4x - 5y D. 5x - 3y = 12 E. - 3x + 7y = -6

OpenStudy (anonymous):

SUBSTITUTE ! :P guess first :)

OpenStudy (anonymous):

I think B!!!!!

OpenStudy (anonymous):

YAY :)

OpenStudy (anonymous):

Wooooo!!!

OpenStudy (anonymous):

party :)

OpenStudy (anonymous):

congratulations!

OpenStudy (anonymous):

Thank you! :)